OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

Use this y-y_1 = m(x-x_1) here slope is -1 by comparing the equation with y = mx+c and comparing m with y = -x+6 you get slope = -1.

OpenStudy (anonymous):

x_1=-4, y_1=2